OpenStudy (anonymous):

4x^3-12x^2 how do i find the 0 roots ?

geerky42 (geerky42):

look like you can factor \(x^2\) out

OpenStudy (anonymous):

4x^2(x-3) okay whats next

geerky42 (geerky42):

good, you factored 4 out too. now according to Zero Product Property, \(4x^2=0\) OR \(x-3=0\) Now can you solve for x?

OpenStudy (anonymous):

3 and 0 ? thats only 2....

OpenStudy (anonymous):

its a cube root.

OpenStudy (anonymous):

should be 3 numbers

geerky42 (geerky42):

yeah, but technically, two of them are same. \(4x^2=0\rightarrow 2x=0~or~2x=0\) So there is only two roots

OpenStudy (anonymous):

thanks

geerky42 (geerky42):

Graph \(4x^3-12x^2\) and you will see that there is only two roots. glad I helped